# Orbitronics in Two-dimensional Materials

### v1

@@ I Introduction
Spintronic phenomena in non-magnetic materials require the presence of spin-orbit coupling (SOC). This relativistic interaction, given by $\lambda\hat{{\bf S}}\cdot\hat{{\bf L}}$, couples the electronic spin operator $\hat{{\bf S}}$ with the orbital angular momentum (OAM) operator $\hat{{\bf L}}$. Its intensity $\lambda$ is material dependent and increases rapidly with the atomic number $Z$ of the constituent elements. A prominent spintronic phenomenon is the spin Hall effect (SHE), in which a transverse spin current is generated in response to a longitudinal applied electric field [5, 6, 7]. In addition to producing pure spin currents for practical applications in spintronic devices, the SHE may serve as a reliable identifier of the quantum spin Hall insulating phase. This topological phase of matter has spurred a considerable amount of research activity in condensed matter physics over the past two decades [8, 9, 10]. However, the need for a strong SOC may greatly limit the options for material selection. For instance, the relatively low atomic number of carbon presents a challenge to manipulating the electronic spin in pristine graphene. To address this limitation, one possible approach involves doping graphene with impurities that exhibit high spin-orbit coupling (SOC) values, or placing it sufficiently close to materials that can induce SOC in graphene through the proximity effect, in order to improve its potential application in spintronics [11, 4]. Similar limitations occur in other materials composed of elements with low atomic numbers [12]. However, materials with naturally high SOC are relatively rare, and their extraction processes often incur significant environmental and economic costs [13].
@@ II Theoretical background
where $\mathcal{J}_{y}^{L_{z}}$ represents the current density component along the $\hat{y}$ direction of the OAM component $L_{z}$, induced by an electric field with intensity $\mathcal{E}_{x}$ applied in the $\hat{x}$ direction. The current density operator is defined by,
@@
Here, the components $\alpha=x,y$ of the velocity operators may be obtained by $\hat{{\rm v}}_{\alpha}({\bf k})=\hbar^{-1}\partial\hat{H}_{\bf k}/\partial{\rm k}_{\alpha}$, $f(\epsilon)$ denotes the usual Fermi-Dirac distribution function and $|u_{n{\bf k}}\rangle$ is the cell-periodic part of the Bloch states with eigenvalue $\epsilon_{n{\bf k}}$.
Fig. 1 shows a schematic representation of the OHE and the resulting accumulation of OAM at the edges of a 2D system due to the transverse orbital current. This effect can occur in the absence of SOC as well as in the presence of spatial inversion and time-reversal symmetry. Within the picture of the mechanism introduced in Ref. [28], SOC partially transfers the transverse orbital current to the spin sector [$\sigma^{L_{z}}_{yx}\rightarrow\sigma^{S_{z}}_{yx}$], giving rise to SHE as a secondary phenomenon resulting from the more fundamental OHE.
The physical picture from Ref. [28] has greatly improved the understanding of OHE. However, it is worth noting that it relies on the intra-atomic approximation and may not provide a complete description compared to other formulations of the OAM operator, which will be discussed later. Additional contributions to this area are detailed in recent reviews focused on the broader aspects of orbitronics [14, 15, 16].
One may generalize the equation above to include other matrix elements of the orbital conductivity tensor $\sigma^{L_{\delta}}_{\mu,\nu}$. This tensor obeys the same symmetry properties as the spin conductivity tensor [58, 59]. However, the physical picture elaborated in Ref. [28] applies to the elements of Eq. (1) and its cyclic permutations in the Cartesian indices $x,y$ and $z$.
Figure 1: Schematic representation of orbital Hall Effect [Eq.(1)]. An electric current passing through the material induces a transverse flow of orbital angular momentum. The color gradient in the rings represents the phase gradient of electron wavefunctions, illustrating how electrons acquire orbital angular momentum and drift in opposite directions on either side of the material.
@@
$\alpha_{\mu\nu}=-e\tau_{p}\frac{\mu_{B}}{\hbar}\sum_{n}\int_{\rm BZ}\frac{d{\bf k}}{(2\pi)^{2}}\left[\frac{\partial f}{\partial\epsilon}\right]_{\epsilon=E_{\rm F}}$
@@
where $\mu_{B}$ denotes the Bohr magneton. Eq. (6) is strongly constrained by crystal symmetry [64]. Since OREE is a Fermi surface effect, the previous equation can be expressed in terms of the charge current as follows:
@@
where $\mathcal{J}_{\nu}$ is the charge current flowing in the $\nu$-direction and $\beta_{\mu\nu}=\alpha_{\mu\nu}/\sigma_{\nu\nu}$, with $\sigma_{\nu\nu}$ being the longitudinal conductivity. Note that while the tensor $\alpha_{\mu\nu}$ is proportional to the momentum scattering time $\tau_{p}$, the tensor $\beta_{\mu\nu}$ is independent of it [$\alpha_{\mu\nu}\propto\tau_{p}^{1}$ and $\beta_{\mu\nu}\propto\tau_{p}^{0}$]. Fig. 2 illustrates the OREE associated with the $\beta_{zx}$ tensor component in Eq. 7.
@@
Figure 2: Schematic representation of orbital Rashba-Edelstein Effect [component $\beta_{zx}$ of Eq.(7)]. An electric current passing through the material induces an orbital magnetization (golden arrows) oriented perpendicular to the current. The circulating electrons, driven by the phase gradient of their wavefunctions (color gradient in the rings), give rise to the orbital angular momentum.
@@
where the first term in Eq. (8) encapsulates the variation of the atomic Coulomb potential due to the motion of the ions, the second term is related to changes in the spin density of the system, and the third term capture the changes in the OAM density. The first mechanism has been recently explored by Han et al. as a way of using the lattice to harvest spin and orbital moments [70]. The second channel connects directly to the spin-orbit torques [71], whereas the third describes the orbital torque in which a non-equilibrium orbital density triggers the magnetization dynamics. This mechanism was proposed by Go and Lee in Ref. [35]. The system they used was a bilayer composed of a non-magnetic material (NM) and a ferromagnetic material (FM) with strong SOC. The NM material is formed by $s$-$p$ hybridized orbitals (similar to the model used in Ref. [28]), while the FM is composed of $d$ orbitals. The equilibrium magnetization ${\bf M}$ of FM interacts with electronic spins density via exchange coupling $H^{\rm FM}_{\rm xc}=\left(J/\hbar\right){\bf M}\cdot\hat{{\bf S}}$. The SOC of FM is given by $H_{\rm so}^{\rm NM}=\left(\alpha^{\rm FM}_{\rm so}/\hbar^{2}\right)\hat{{\bf L}}\cdot\hat{{\bf S}}$. When an electric field is applied, a transverse orbital current flows in the NM material due to the OHE. This orbital current is then transferred to the FM, injecting an OAM density, and the angular momentum of this nonequilibrium distribution is transferred to the spins via SOC. This process induces a torque in the magnetization of the FM material. The torque is given by
@@
where $\langle\hat{{\bf S}}\rangle^{\rm FM}$ is the spin accumulation produced in FM. Note that although the expression of Eq. (9) for OT depends on spin accumulation, it results from the injection of the OAM due to OHE in the NM material described by the third term in (8). This expression can be decomposed into field-like and damping-like components. While the field-like component arises from the intra-band term of linear response theory, the damping-like term is associated with inter-band contribution [35]. This phenomenon has been experimentally analyzed in various NM/FM bilayers and represents an important route for manipulating the OAM degree of freedom [72, 73, 74]. The reciprocal effect, orbital pumping, has been studied theoretically and its experimental evidence has also been reported [39, 75, 76, 70, 77].
@@
The effects described in Eqs.(1-9), along with their reciprocal effects, can be viewed as building blocks for orbitronics devices within the generation-manipulation-detection paradigm, which is aimed at enabling efficient storage and processing of information using electronic OAM. Nevertheless, this has not exhausted the pool of orbitronic phenomena that could be useful for this task. Distinct mechanisms of orbital-charge conversion [78], and the interaction between OAM and pulsed light signals [79, 80, 81] are other topics currently under theoretical and experimental investigation. Here, we focused on the three building blocks, whose mechanisms and key references were outlined above.
@@ III Orbital angular momentum in 2D materials
The mechanism introduced in Ref. [28] to explain the OHE depends on the presence of orbital textures, $\bf{k}$-dependent orbital states, even in systems with inversion and time-reversal symmetry[82]. These textures, which describe the spatial distribution of orbital character, are an important feature in many 2D materials. Within the framework of the intra-atomic approximation (see the next subsection), these textures are associated with $\bf{k}$-dependent hopping processes involving different orbitals. However, 2D materials with broken inversion symmetry can give rise to OAM textures, described by $\bm{\ell}_{n}({\bf k})=\bra{u_{n{\bf k}}}\hat{{\bf L}}\ket{u_{n{\bf k}}}$, where |$\ket{u_{n{\bf k}}}$ represents the periodic part of the Bloch wave functions. Such OAM textures highlight the unique role of symmetry breaking in 2D materials. They are found in various multi-orbital solids, including TMDs [45], bismuthene [46], and borophene [83].
@@
where, $\vec{\nabla}_{\bf k}=\partial_{k_{x}}\vec{x}+\partial_{k_{y}}\vec{y}$, and |$\ket{u_{n(m){\bf k}}}$ is the periodic part of Bloch wave-function for the band with energy $\epsilon_{n(m){\bf k}}$. Eq. (10) is analogous to the expression obtained from the semi-classical approach but is not restricted to the nearly degenerate subspace. The same expression was obtained by Ref. [93]. As pointed out in Ref. [94], an additional term must be added to Eq. (10) to ensure gauge invariance in the case of spatial energy bands. Recently, an additional quantum term has been predicted [95].
@@
that was introduced by Canonico et al. [102]. Within this approach, the off-diagonal elements of the position operator $\bm{r}$ are computed from the perturbation theory methods as $\langle i|r_{\alpha}|j\rangle=i\hbar\frac{\langle i|\hat{{\rm v}}_{\alpha}|j\rangle}{E_{j}-E_{i}}$ ($\alpha=x,y,z$), where $|i\rangle$ and $E_{i}$ are the $i$-th energy eigenstate of Hamiltonian $\hat{H}$ with eigenvalue $E_{i}$. $G^{-}(\hat{H},E)$ and $G^{+}(\hat{H},E)$ are the advanced and retarded Green’s functions respectively and $\varepsilon_{\alpha\beta\gamma}$ represent the Levi-Civita symbol. The formulation in Eq. (12) has been used to calculate the OHE due to the OMM in real space for disordered systems [102], where the expressions in Eqs. (10), (11) and (12) are used to describe the transport of OAM via the orbital current definition shown in Eq. (2).
@@
The first line of Eq. (13) represents the macroscopic orbital magnetization as an integral of a local quantity [24, 26]. $A$ is the area of the sample, $\mu$ is the chemical potential, $\mathcal{P}=\sum_{\epsilon_{n}<\mu}|\phi_{n}\rangle\langle\phi_{n}|$ is the projector over the occupied states, and $|\hat{H}-\mu\mathbb{1}|=\big{(}\hat{H}-\mu\mathbb{1}\big{)}\big{(}\mathbb{1}-2\mathcal{P}\big{)}$ [103]. This expression exploits the spectral representation of the orbital moment due to its independence on the characteristics of the energy states and is suitable for computing the orbital moment in disordered materials. The second line in Eq. (13) in turn, is usual expression that exploits the periodicity of clean systems and represents it as a bulk property [25, 24], i.e., an integral of a geometric quantity over the BZ. Note that in both representations, the orbital magnetization does not depend on the explicit representation of the multi-orbital nature of the band structure. It depends only on the eigenstates and eigenvalues of the Hamiltonian of magnetic material.

### v2

@@ I Introduction
Spin current is a key concept in the field of spintronics. The generation of pure spin current in time-reversal symmetric systems requires the presence of spin-orbit coupling (SOC). This relativistic interaction, given by $\lambda\hat{{\bf S}}\cdot\hat{{\bf L}}$, couples the electronic spin operator $\hat{{\bf S}}$ with the orbital angular momentum (OAM) operator $\hat{{\bf L}}$. Its intensity $\lambda$ is material dependent and increases rapidly with the atomic number $Z$ of the constituent elements. A prominent spintronic phenomenon is the spin Hall effect (SHE), in which a transverse spin current is generated in response to a longitudinal applied electric field [5, 6, 7]. In addition to producing pure spin currents for practical applications in spintronic devices, the SHE may serve as a reliable identifier of the quantum spin Hall insulating phase. This topological phase of matter has spurred a considerable amount of research activity in condensed matter physics over the past two decades [8, 9, 10]. However, the need for a strong SOC may greatly limit the options for material selection. For instance, the relatively low atomic number of carbon presents a challenge to manipulating the electronic spin in pristine graphene. To address this limitation, one possible approach involves doping graphene with impurities that exhibit high spin-orbit coupling (SOC) values, or placing it sufficiently close to materials that can induce SOC in graphene through the proximity effect, in order to improve its potential application in spintronics [11, 4]. Similar limitations occur in other materials composed of elements with low atomic numbers [12]. However, materials with naturally high SOC are relatively rare, and their extraction processes often incur significant environmental and economic costs [13].
@@ II Theoretical background
where $\mathcal{J}_{y}^{L_{z}}$ represents the current density component along the $\hat{y}$ direction of the OAM component $L_{z}$, induced by an electric field with intensity $\mathcal{E}_{x}$ applied in the $\hat{x}$ direction.
Fig. 1 shows a schematic representation of the OHE and the resulting accumulation of OAM at the edges of a 2D system due to the transverse orbital current. It is worth mentioning that the orbital current is a quantum mechanical expectation value of the OAM flux carried by the conduction electrons; it does not represent a transverse flow of charge carriers. Rather, it describes how electrons, while moving longitudinally under an electric field, acquire OAM, resulting in a net angular momentum flow without any accompanying net charge transport in the transverse direction.
This effect can occur in the absence of SOC as well as in the presence of spatial inversion and time-reversal symmetry. Within the picture of the mechanism introduced in Ref. [28], SOC partially transfers the transverse orbital current to the spin sector [$\sigma^{L_{z}}_{yx}\rightarrow\sigma^{S_{z}}_{yx}$], giving rise to SHE as a secondary phenomenon resulting from the more fundamental OHE. Other theoretical aspects of the OHE are detailed in recent reviews [14, 15, 16].
One may generalize equation 1 to include other matrix elements of the orbital conductivity tensor $\sigma^{L_{\delta}}_{\mu,\nu}$. This tensor obeys the same symmetry properties as the spin conductivity tensor [58, 59]. However, the physical picture elaborated in Ref. [28] applies to the elements of Eq. (1) and its cyclic permutations in the Cartesian indices $x,y$ and $z$.
The current density operator is defined by,
@@
Figure 1: OHE in a 2D system. Schematic representation of the orbital Hall effect [Eq.(1)]. A longitudinal charge current (black arrow) driven by an external electric field (red arrow) leads to a transverse flow of orbital angular momentum (OAM). This orbital current results in the accumulation of opposite OAM at the lateral edges of the sample, represented by golden arrows and phase rings. The color gradient in the rings indicates the phase structure of the electron wavefunctions associated with nonzero OAM. The splitting of the charge current line is purely pictorial, used to illustrate the opposite deflection of electrons carrying different signs of OAM. No transverse charge current is present.
@@
Here, the components $\alpha=x,y$ of the velocity operators may be obtained by $\hat{{\rm v}}_{\alpha}({\bf k})=\hbar^{-1}\partial\hat{H}_{\bf k}/\partial{\rm k}_{\alpha}$, $f(\epsilon)$ denotes the usual Fermi-Dirac distribution function and $|u_{n{\bf k}}\rangle$ is the cell-periodic part of the Bloch states with eigenvalue $\epsilon_{n{\bf k}}$. Figure 2 illustrates the valence-band orbital Berry curvature for a monolayer of 2H-TMD, a prototypical multi-orbital Dirac material discussed in this review. In solids, the Berry curvature plays a role analogous to a magnetic field in momentum space, encoding how electronic wavefunctions twist or acquire geometric phase as the crystal momentum changes [60, 61]. The orbital Berry curvature refines this concept by taking into account the orbital character of the wavefunctions—effectively weighting the conventional Berry curvature by the expectation value of the OAM operator [28]. It captures how electrons with specific orbital identities respond to external fields, and directly contributes to orbital transport phenomena such as the OHE. While the conventional Berry curvature of monolayer 2H-TMD valence bands has opposite signs at the ${\bf K}$ and ${\bf K}^{\prime}$ valleys—leading to effects like the valley Hall effect—the orbital Berry curvature, as shown in Fig. 2 (d), has the same sign at both valleys. This distinction is crucial, as it gives rise to a net orbital Hall conductivity even in centrosymmetric systems that preserve time-reversal symmetry [48, 49], and is responsible for the emergence of the orbital Hall insulating phase discussed in later sections [45].
Figure 2: Berry and Orbital Berry Curvature in 2H-TMD Monolayers. (a) Schemmatic illustration of the calculated energy bands for non-centrosymmetric 2H-TMD monolayer (a typical massive Dirac material) around the bottom of its conduction band and the top of the valence bands located at ${\bf K}$ and ${\bf K^{\prime}}$ valleys of Brillouin zone (BZ). The black curve shows the spectrum without considering spin-orbit coupling (SOC). The dashed brown ($s=\uparrow$) and dotted purple ($s=\downarrow$) curves show the spin splitting of the valence bands caused by SOC. (b) The valence band Berry curvature near the valleys of the BZ. (c) The z-component of orbital angular momentum (OAM) near the valleys of the BZ. (d) The valence band orbital Berry curvature near the valleys of the BZ for a 2H-TMD monolayer. In the panels (c) and (d), the blue curves represent the results in the intra-atomic approximation, while the red curves show the results for OAM originating from the magnetic moment of Bloch states [49]. In panels (b)–(d), the SOC of the TMD is not considered, and a spin-degeneracy factor is included in the results for the Berry and orbital Berry curvatures. The low-energy theory of 2H-TMD monolayer with parameters for MoS2 is used in these results.
@@
$\alpha_{\mu\nu}=-e\tau_{p}\frac{\mu_{B}}{\hbar}\sum_{n}\int_{\rm BZ}\frac{d{\bf k}}{(2\pi)^{2}}\left[\frac{\partial f}{\partial\epsilon}\right]_{\epsilon=\epsilon_{n{\bf k}}}$
@@
where $\mu_{B}$ denotes the Bohr magneton and $\tau_{p}$ represents the momentum relaxation time. Eq. (6) is strongly constrained by crystal symmetry [66]. Since OREE is a Fermi surface effect, the previous equation can be expressed in terms of the charge current as follows:
@@
where $\mathcal{J}_{\nu}$ is the charge current flowing in the $\nu$-direction and $\beta_{\mu\nu}=\alpha_{\mu\nu}/\sigma_{\nu\nu}$, with $\sigma_{\nu\nu}$ being the longitudinal conductivity. Note that since $\alpha_{\mu\nu}\propto\tau_{p}$, the tensor $\beta_{\mu\nu}$ is independent of $\tau_{p}$. Fig. 3 illustrates the OREE associated with the $\beta_{zx}$ tensor component in Eq. 7.
@@
Figure 3: OREE in a 2D system. Schematic representation of orbital Rashba-Edelstein Effect [component $\beta_{zx}$ of Eq.(7)]. An electric current passing through the material induces an orbital magnetization (golden arrows) oriented perpendicular to the current. The circulating electrons, driven by the phase gradient of their wavefunctions (color gradient in the rings), give rise to the orbital angular momentum.
@@
where the first term in Eq. (8) encapsulates the variation of the atomic Coulomb potential due to the motion of the ions, the second term is related to changes in the spin density of the system, and the third term capture the changes in the OAM density. The first mechanism has been recently explored by Han et al. [72] as a way of using the lattice to harvest spin and orbital moments. The second channel connects directly to the spin-orbit torques [73], whereas the third describes the orbital torque in which a non-equilibrium orbital density triggers the magnetization dynamics. This mechanism was proposed by Go and Lee in Ref. [35]. The system they used was a bilayer comprising a non-magnetic material (NM) and a ferromagnetic material (FM) with strong SOC. The NM material is formed by $s$-$p$ hybridized orbitals (similar to the model used in Ref. [28]), while the FM includes $d$ orbitals. The equilibrium magnetization ${\bf M}$ of FM interacts with electronic spins density via exchange coupling $H^{\rm FM}_{\rm xc}=\left(J/\hbar\right){\bf M}\cdot\hat{{\bf S}}$. The SOC of FM is given by $H_{\rm so}^{\rm FM}=\left(\alpha^{\rm FM}_{\rm so}/\hbar^{2}\right)\hat{{\bf L}}\cdot\hat{{\bf S}}$. When an electric field is applied, a transverse orbital current flows in the NM material due to the OHE. This orbital current is then transferred to the FM, injecting an OAM density, and the angular momentum of this nonequilibrium distribution is transferred to the spins via SOC. This process induces a torque in the magnetization of the FM material. The torque is given by
@@
where $\langle\hat{{\bf S}}\rangle^{\rm FM}$ is the spin accumulation produced in FM. Note that although the expression of Eq. (9) for OT depends on spin accumulation, it results from the injection of the OAM due to OHE in the NM material described by the third term in (8). This expression can be decomposed into field-like and damping-like components. While the field-like component arises from the intra-band term of linear response theory, the damping-like term is associated with inter-band contribution [35]. This phenomenon has been experimentally analyzed in various NM/FM bilayers and represents an important route for manipulating the OAM degree of freedom [74, 75, 76]. The reciprocal effect, orbital pumping, has been studied theoretically and its experimental evidence has also been reported [39, 77, 78, 72, 79]. It is worth mentioning that the model described above assumes that the magnetization of a ferromagnet is primarily attributed to the spin sector. In the cases where the ferromagnet has significant contributions from the orbital sector to its magnetization, the model can be generalized [80].
@@
The effects described in Eqs.(1-9), along with their reciprocal effects, can be viewed as building blocks for orbitronics devices within the generation-manipulation-detection paradigm, which is aimed at enabling efficient storage and processing of information using electronic OAM. Nevertheless, this has not exhausted the pool of orbitronic phenomena that could be useful for this task. Distinct mechanisms of orbital-charge conversion [81], orbital Ruderman-Kittel-Kasuya-Yosida interaction [80], and the interaction between OAM and pulsed light signals [82, 83, 84] are other topics currently under theoretical and experimental investigation. Here, we focused on the three building blocks, whose mechanisms and key references were outlined above.
@@ III Orbital angular momentum in 2D materials
The mechanism introduced in Ref. [28] to explain the OHE depends on the presence of orbital textures, $\bf{k}$-dependent orbital states, even in systems with inversion and time-reversal symmetry[85]. These textures, which describe the spatial distribution of orbital character, are an important feature in many 2D materials. Within the framework of the intra-atomic approximation (see the next subsection), these textures are associated with $\bf{k}$-dependent hopping processes involving different orbitals. However, 2D materials with broken inversion symmetry can give rise to OAM textures, described by $\bm{\ell}_{n}({\bf k})=\bra{u_{n{\bf k}}}\hat{{\bf L}}\ket{u_{n{\bf k}}}$, where |$\ket{u_{n{\bf k}}}$ represents the periodic part of the Bloch wave functions. Such OAM textures highlight the unique role of symmetry breaking in 2D materials. They are found in various multi-orbital solids, including monolayers of 2H-TMDs [45], bismuthene [46], and borophene [86].
@@
where, $\vec{\nabla}_{\bf k}=\partial_{k_{x}}\vec{x}+\partial_{k_{y}}\vec{y}$, and |$\ket{u_{n(m){\bf k}}}$ is the periodic part of Bloch wave-function for the band with energy $\epsilon_{n(m){\bf k}}$. Eq. (10) is analogous to the expression obtained from the semi-classical approach but is not restricted to the nearly degenerate subspace. The same expression was obtained by Ref. [96]. As pointed out in Ref. [97], an additional term must be added to Eq. (10) to ensure gauge invariance in the case of spaced energy bands. Recently, an additional quantum term has been predicted [98].
@@
that was introduced by Canonico et al. [105]. Within this approach, the off-diagonal elements of the position operator $\bm{r}$ are computed from the perturbation theory methods as $\langle i|r_{\alpha}|j\rangle=i\hbar\frac{\langle i|\hat{{\rm v}}_{\alpha}|j\rangle}{E_{j}-E_{i}}$ ($\alpha=x,y,z$), where $|i\rangle$ and $E_{i}$ are the $i$-th energy eigenstate of Hamiltonian $\hat{H}$ with eigenvalue $E_{i}$. $G^{-}(\hat{H},E)$ and $G^{+}(\hat{H},E)$ are the advanced and retarded Green’s functions respectively and $\varepsilon_{\alpha\beta\gamma}$ represent the Levi-Civita symbol. The formulation in Eq. (12) has been used to calculate the OHE due to the OMM in real space for disordered systems [105]. The expressions in Eqs. (10), (11) and (12) are used to describe the transport of OAM via the orbital current definition shown in Eq. (2).
@@
The first line of Eq. (13) represents the macroscopic orbital magnetization as an integral of a local quantity [24, 26]. $A$ is the area of the sample, $\mu$ is the chemical potential, $\mathcal{P}=\sum_{\epsilon_{n}<\mu}|\phi_{n}\rangle\langle\phi_{n}|$ is the projector over the occupied states, and $|\hat{H}-\mu\mathbb{1}|=\big{(}\hat{H}-\mu\mathbb{1}\big{)}\big{(}\mathbb{1}-2\mathcal{P}\big{)}$ [106]. This expression exploits the spectral representation of the orbital moment due to its independence on the characteristics of the energy states and is suitable for computing the orbital moment in disordered materials. The second line in Eq. (13) in turn, is the usual expression that exploits the periodicity of clean systems and represents it as a bulk property [25, 24], i.e., an integral of a geometric quantity over the BZ. Note that in both equations, the orbital magnetization does not depend on the explicit representation of the multi-orbital nature of the band structure. It depends only on the eigenstates and eigenvalues of the Hamiltonian of magnetic material.